# AEDR: Training-Free AI-Generated Image Attribution via Autoencoder Double-Reconstruction

### v1

@@ Abstract
The rapid advancement of image-generation technologies has made it possible for anyone to create photorealistic images using generative models, raising significant security concerns. To mitigate malicious use, tracing the origin of such images is essential. Reconstruction-based attribution methods offer a promising solution, but they often suffer from reduced accuracy and high computational costs when applied to state-of-the-art (SOTA) models. To address these challenges, we propose AEDR (AutoEncoder Double-Reconstruction), a novel training-free attribution method designed for generative models with continuous autoencoders. Unlike existing reconstruction-based approaches that rely on the value of a single reconstruction loss, AEDR performs two consecutive reconstructions using the model’s autoencoder, and adopts the ratio of these two reconstruction losses as the attribution signal. This signal is further calibrated using the image homogeneity metric to improve accuracy, which inherently cancels out absolute biases caused by image complexity, with autoencoder-based reconstruction ensuring superior computational efficiency. Experiments on eight top latent diffusion models show that AEDR achieves 25.5% higher attribution accuracy than existing reconstruction-based methods, with requiring only 1% of the computational time.
@@ 1 Introduction
Latent diffusion models Labs (2024); Razzhigaev et al. (2023); Tang et al. (2022) have emerged as the dominant paradigm in image generation due to their exceptional capabilities in producing high-resolution and photorealistic content. Their effectiveness in modeling complex data distributions has facilitated numerous applications across diverse fields, such as digital art, advertising, and virtual reality Dhariwal and Nichol (2021). Recent models such as Stable Diffusion 3.5 Esser et al. (2024) enable rapid and controllable image synthesis, supporting a wide range of real-world applications.
While the powerful capabilities of AI-generated images are widely embraced, concerns about potential misuse are growing Carlini et al. (2023); Chen et al. (2020); Ong et al. (2021); Zhao et al. (2021). For instance, unscrupulous vendors may repurpose outputs from third-party models as their own, falsely promoting model performance and thereby misleading consumers while undermining fair market competition. Furthermore, malicious actors can present outputs from commercial models Betker et al. (2023) as original creations to gain reputational and financial advantages, flagrantly violating the intellectual property rights of model developers Li et al. (2024). Reliable image origin attribution, therefore, has become indispensable for accurately identifying the responsible entities behind such generated content.
Existing image attribution methods generally fall into three primary categories. Watermark-based methods Luo et al. (2009); Pereira and Pun (2000); Swanson et al. (1996); Wen et al. (2023); Yang et al. (2024, 2025) embed invisible or semi-visible marks during image generation, enabling attribution through watermark detection. Fingerprint-based methods Ding et al. (2021); Yu et al. (2019, 2020, 2021) introduce model-specific signatures during training or by modifying the architecture, and rely on supervised classifiers to detect these fingerprints in generated images. However, both methods require additional operations during training or inference, which may degrade the visual quality of the generated images. In contrast, passive detection approaches, which do not require modifications to the training process of target models, are more practical and acceptable in real-world applications. Among these, reconstruction-based methods Wang et al. (2023, 2024) have achieved excellent performance. They use gradient information from the generative model to reconstruct the input image, using the reconstruction loss between the original and reconstructed images serving as the attribution signal. These methods do not modify model parameters or the generation pipeline, and thus preserve generative performance, making them one of the most promising approaches for image attribution.
(a) Gradient-based reconstruction using SD2base.
(b) Gradient-based reconstruction using FLUX.
Figure 1: Gradient-based reconstruction methods exhibit different loss distributions.
@@
Figure 3: Double-Reconstruction loss variation and loss ratio.
@@
Existing gradient-based reconstruction methods typically follow this observation: images that belong to the target model generally exhibit lower reconstruction losses than non-belonging images (see Figure 1(a)). Thus, the two types of images can be distinguished based on a single reconstruction loss value. However, with increasingly powerful generative models such as FLUX Labs (2024), these methods Wang et al. (2023, 2024) tend to produce extremely low reconstruction losses for both belonging and non-belonging images, causing significant overlap in the loss distributions and thereby reducing attribution accuracy (see Figure 1(b)). Moreover, the gradient-guided process is computationally expensive and complex, making these methods impractical for real-world applications.
@@
To address these limitations, we prefer to adopt autoencoder-based reconstruction as the foundation for an efficient attribution. However, using reconstruction loss directly is susceptible to the inherent complexity of the input image Choi et al. (2024); Ricker et al. (2024): Images with simple textures naturally have lower reconstruction loss, while those with complex textures correspondingly have higher reconstruction loss. To mitigate this bias introduced by image complexity, we propose an attribution method based on the loss ratio across double-reconstructions by the autoencoder. The key insight is that the autoencoder has learned to capture representative features of samples from the training distribution. As illustrated in Figure 3(a) and Figure 3(b), we observe that when the autoencoder reconstructs a belonging image twice, the resulting reconstruction losses are nearly identical, since the image lies within the model’s distribution. In contrast, for the non-belonging image, which initially lies outside the model distribution, the first reconstruction effectively projects it into the distribution, leading to a noticeably lower loss in the second reconstruction. Consequently, if we compute the ratio of the first to the second reconstruction loss, the value tends to be close to 1 for belonging images, but significantly greater than 1 for non-belonging ones (see Figure 3(c)).
@@
Motivated by this insight, we propose an image attribution method based on AutoEncoder Double-Reconstruction, namely AEDR. Unlike existing gradient-guided approaches that utilize single reconstruction loss, AEDR employs the loss ratio from double reconstructions, calibrated by an image homogeneity metric. This approach mitigates discrepancies arising from image texture complexity. Additionally, we implement a Kernel Density Estimation (KDE) method Kim and Scott (2012) for adaptive threshold selection, which assumes no prior knowledge of underlying data distributions, thereby enhancing the adaptability across diverse generative models. Our contributions are summarized as follows:
@@
  • We reveal that existing reconstruction-based generated image attribution methods, which rely on the single reconstruction loss, face significant challenges when applied to SOTA generative models. We identify a key discrepancy: under autoencoder-based double-reconstruction, belonging images show significantly higher latent feature consistency across their reconstructions than non-belonging ones.
@@
  • We propose AEDR, an attribution method that leverages the ratio of losses from two successive reconstructions and calibrates it based on image homogeneity. AEDR performs direct reconstruction using the model’s autoencoder, eliminating the need for gradient-based optimization or additional training.
@@
  • Extensive experiments on eight state-of-the-art or widely used generative models demonstrate that AEDR improves the attribution accuracy by an average of 25.5% and reduces the attribution time to just 1% of the best-performing baseline.
@@ 2 Related Work
Early frameworks like Generative Adversarial Networks (GANs) Goodfellow et al. (2014), Variational Autoencoders (VAEs) Kingma and Welling (2014), and autoregressive methods such as PixelCNN Van Den Oord et al. (2016) laid the groundwork for realistic image synthesis. However, diffusion models have recently become the leading approach. Denoising Diffusion Probabilistic Models (DDPMs) Ho et al. (2020) achieve remarkable image fidelity by iteratively adding and reversing Gaussian noise, using simplified objectives and U-Net architectures. Building on this, Latent Diffusion Models (LDMs) Rombach et al. (2022) improved efficiency by processing diffusion within a compact latent space, enabling rapid, high-resolution image generation with significantly lower computational requirements. Many SOTA image generative models, such as Stable Diffusion Esser et al. (2024); Podell et al. (2023) and FLUX Labs (2024), also follow this paradigm.
@@
As generative images grow increasingly photorealistic, the need for reliable detection is increasing Chen et al. (2023). Early work by Marra combined CycleGAN with steganalysis to identify GAN-generated images, laying a foundation for this field. Most existing methods frame the task as a binary classification problem, distinguishing between real and generated images. These approaches often rely on discriminative texture features Chen and Yashtini (2024); Gye et al. (2025); Konstantinidou et al. (2025); Li et al. (2025); Raza et al. (2024) or frequency domain signals Corvi et al. (2023); Dong et al. (2022); Dzanic et al. (2020); Frank et al. (2020); Jeong et al. (2022) as core detection cues. Recent approaches such as AEROBLADE Ricker et al. (2024) and HFI Choi et al. (2024) adopt reconstruction losses via autoencoders and use LPIPS as a detection metric. Although these methods demonstrate promising performance in generative image detection, they do not address the more challenging task of origin attribution, which is the focus of this paper.
@@
Current methods for attributing the source of the generated images can be broadly classified into three types: watermark-based methods Luo et al. (2009); Pereira and Pun (2000); Swanson et al. (1996); Wen et al. (2023); Yang et al. (2024, 2025), fingerprint-based methods Ding et al. (2021); Yu et al. (2019, 2020, 2021), and reconstruction-based methods Wang et al. (2023, 2024). Watermark-based methods embed model-specific information into the generated image and retrieve this information during detection to identify the source. Fingerprint-based methods inject unique patterns into the model during training and rely on dedicated classifiers for attribution. However, both approaches require additional intervention during image generation or model training and may compromise the quality of the generated images. In contrast, reconstruction-based methods enable origin attribution without the need for any extra operations. For example, both RONAN Wang et al. (2023) and LatentTracer Wang et al. (2024) adopt gradient-based reconstruction approaches, using the reconstruction loss between the reconstructed and original images as the attribution signal. These methods have been shown to be effective for certain generative models; however, their attribution accuracy degrades significantly as model performance increases.
3 Problem Formulation
In this section, we formally define the key problems related in this work as follows:
Definition 1 (Images Attributed to the Target Model). Given an image generative model $\mathcal{M}:\mathcal{I}\rightarrow\mathcal{X}_{\mathcal{M}}$ , where $\mathcal{I}$ is the input space and $\mathcal{X}_{\mathcal{M}}\subset\mathcal{X}$ denotes the subset of images produced by $\mathcal{M}$, with $\mathcal{X}$ representing the full image space. A test image $x$ is classified as a belonging image if and only if $x\in\mathcal{X}_{\mathcal{M}}$; if $x\notin\mathcal{X}_{\mathcal{M}}$, it is classified as a non-belonging image.
Definition 2 (Autoencoder-Based Reconstruction). The autoencoder $\mathcal{R}$ of the target model $\mathcal{M}$ comprises an encoder $\mathcal{E}:\mathcal{X}\rightarrow\mathcal{Z}$ and a decoder $\mathcal{D}:\mathcal{Z}\rightarrow\mathcal{X}$, where $\mathcal{X}$ denotes the pixel space and $\mathcal{Z}$ denotes the latent space. The autoencoder-based reconstruction process is defined as $x^{*}=\mathcal{R}(x)=\mathcal{D}(z)=\mathcal{D}(\mathcal{E}(x))$, where $x\in\mathcal{X}$ is the input image, $z=\mathcal{E}(x)$ is its latent representation, and $x^{*}=\mathcal{D}(z)$ is the reconstructed image.
Definition 3 (Double-Reconstruction Based Origin Attribution). Given a test image $x$ and a generative model $\mathcal{M}$ with autoencoder $\mathcal{R}$, two rounds of encoding-decoding yield reconstructions $x^{*}=\mathcal{D}(\mathcal{E}(x))$ and $x^{**}=\mathcal{D}(\mathcal{E}(x^{*}))$. The attribution signal $t^{\prime}$ is defined as the ratio of the reconstruction losses $\mathcal{L}(x^{*},x)$ and $\mathcal{L}(x^{**},x^{*})$, further calibrated by an image homogeneity metric. If and only if $t^{\prime}<\tau$, where $\tau$ is a predefined threshold, $x$ is classified as a belonging image; otherwise, it is considered a non-belonging image.
@@ 4 Method
This section presents a detailed overview of the proposed method, AEDR. The process begins by reconstructing the input image twice using an autoencoder, from which the reconstruction loss ratio is computed. To enhance robustness, this ratio is calibrated with an image homogeneity metric, resulting in the generation of the attribution signal. Finally, the attribution signal is compared against a precomputed threshold to determine the final attribution outcome. The overall AEDR framework is illustrated in Figure 4.
4.1 Goal and Capability of the Detector
Goal. The detector aims to enable fast and accurate origin attribution for AI-generated images without modifying the training or inference of the target model $\mathcal{M}$, and without applying any post-processing to its outputs. Formally, the attribution is defined as a binary function $F:\{\mathcal{M},x\}\rightarrow\{0,1\}$, where $F(\mathcal{M},x)=0$ denotes that image $x$ is attributed to model $\mathcal{M}$ (belonging image), and $F(\mathcal{M},x)=1$ indicates that it is not (non-belonging image).
Capability. The detector has access only to the autoencoder $\mathcal{R}$ of the target model $\mathcal{M}$, enabling reconstruction via encoder and decoder queries. In contrast to methods like RONAN Wang et al. (2023) and LatentTracer Wang et al. (2024), which require full white-box access and gradient information, AEDR operates exclusively through autoencoder interaction, making it more suitable for scenarios with strict security constraints or limited computational resources.
4.2 Double-Reconstruction Based on Autoencoder
The specific workflow is as follows: Given the autoencoder $\mathcal{R}$ of the target model $\mathcal{M}$ and a test image $x$, we perform the following two reconstruction steps.
First Reconstruction:
@@ 4.2 Double-Reconstruction Based on Autoencoder
Here, $x^{*}$ and $x^{**}$ denote the reconstructed images obtained from the first and second passes through the autoencoder $\mathcal{R}$. The corresponding reconstruction losses are denoted by $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, both computed using mean squared error (MSE). The advantage of using MSE is demonstrated in Experiment 5.5. AEDR focuses on analyzing the trend of variation in reconstruction losses for the test image $x$ during the double-reconstruction process. Specifically, we define an uncalibrated attribution signal $t$ as the ratio between the first and second reconstruction losses, formally given by:

### v2

@@ Abstract
The rapid advancement of image-generation technologies has made it possible for anyone to create photorealistic images using generative models, raising significant security concerns. To mitigate malicious use, tracing the origin of such images is essential. Reconstruction-based attribution methods offer a promising solution, but they often suffer from reduced accuracy and high computational costs when applied to state‑of‑the‑art (SOTA) models. To address these challenges, we propose AEDR (AutoEncoder Double-Reconstruction), a novel training‑free attribution method designed for generative models with continuous autoencoders. Unlike existing reconstruction‑based approaches that rely on the value of a single reconstruction loss, AEDR performs two consecutive reconstructions using the model’s autoencoder, and adopts the ratio of these two reconstruction losses as the attribution signal. This signal is further calibrated using the image homogeneity metric to improve accuracy, which inherently cancels out absolute biases caused by image complexity, with autoencoder‑based reconstruction ensuring superior computational efficiency. Experiments on eight top latent diffusion models show that AEDR achieves 25.5% higher attribution accuracy than existing reconstruction‑based methods, while requiring only 1% of the computational time.
Codehttps://github.com/wangchao0708/AEDR
@@ 1 Introduction
Latent diffusion models (Podell et al. 2023; Batifol et al. 2025; Rombach et al. 2022; Razzhigaev et al. 2023) have emerged as the dominant paradigm in image generation due to their exceptional capabilities in producing high-resolution and photorealistic content. Their effectiveness in modeling complex data distributions has facilitated numerous applications across diverse fields, such as digital art, advertising, and virtual reality (Dhariwal and Nichol 2021). Recent models such as Stable Diffusion 3.5 (Esser et al. 2024) and FLUX.1-dev enable rapid and controllable image synthesis, supporting a wide range of real-world applications.
While the powerful capabilities of AI‑generated images are widely embraced, concerns about potential misuse are growing (Carlini et al. 2023; Chen et al. 2020; Ong et al. 2021; Zhao et al. 2021). For instance, unscrupulous vendors may repurpose outputs from third-party models as their own, falsely promoting model performance and thereby misleading consumers while undermining fair market competition. Furthermore, malicious actors can present outputs from commercial models (Betker et al. 2023) as original creations to gain reputational and financial advantages, flagrantly violating the intellectual property rights of model developers (Li et al. 2024). Therefore, reliable image origin attribution has become indispensable for accurately identifying the responsible entities behind such generated content.
(a) Gradient-based reconstruction using SD2base $(\sim 10^{-4})$.
(b) Gradient-based reconstruction using FLUX $(\sim 10^{-5})$.
Figure 1: Gradient-based reconstruction methods exhibit different loss distributions. These differences lead to attribution failures on the latest model, such as FLUX.
@@
Figure 2: Autoencoder-based double-reconstruction loss variation and loss ratio.
Existing image attribution methods generally fall into three primary categories. Watermark-based methods (Luo et al. 2009; Swanson et al. 1996; Wen et al. 2023; Yang et al. 2024, 2025) embed invisible or semi-visible marks during image generation, enabling attribution through watermark detection. Fingerprint-based methods (Ding et al. 2021; Yu et al. 2020, 2021) introduce model-specific signatures during training or by modifying the architecture, and rely on supervised classifiers to detect these fingerprints in generated images. However, some of these methods may degrade the visual quality of the generated images due to additional operations during training or inference. In contrast, passive detection approaches, which do not require modifications to the training process of target models, are more practical and acceptable in real-world applications. Among these, reconstruction-based methods (Wang et al. 2023, 2024) have achieved excellent performance. They use gradient information from the generative model to reconstruct the input image, with the reconstruction loss between the original and reconstructed images serving as the attribution signal. These methods do not modify model parameters or the generation pipeline, and thus preserve generative performance, making them one of the most promising approaches for image attribution.
@@
Existing gradient-based reconstruction methods typically follow this observation: images that belong to the target model generally exhibit lower reconstruction losses than non-belonging images (see Figure 1(a)). Thus, the two types of images can be distinguished based on a single reconstruction loss value. However, with increasingly powerful generative models such as FLUX (Batifol et al. 2025), these methods (Wang et al. 2023, 2024) tend to produce extremely low reconstruction losses for both belonging and non-belonging images, causing significant overlap in the loss distributions and thereby reducing attribution accuracy (see Figure 1(b)). Moreover, the gradient-guided reconstruction process is computationally expensive and complex, making these methods impractical for real-world applications.
@@
To address these limitations, we prefer to adopt autoencoder-based reconstruction as the foundation for an efficient attribution. However, using reconstruction loss directly is susceptible to the inherent complexity of the input image (Choi et al. 2024; Ricker et al. 2024): Images with simple textures naturally have lower reconstruction loss, while those with complex textures correspondingly have higher reconstruction loss (Further details are provided in Sections 2 and 3 of the Technical Appendix). To mitigate this bias introduced by image complexity, we propose an attribution method based on the loss ratio across double-reconstructions by the autoencoder. The key insight is that the autoencoder has learned to capture representative features of samples from the training distribution. As illustrated in Figure 2(a) and Figure 2(b), we observe that when the autoencoder reconstructs a belonging image twice, the resulting reconstruction losses are nearly identical, since the image lies within the model’s distribution. In contrast, for the non-belonging image, which initially lies outside the model distribution, the first reconstruction effectively projects it into the distribution, leading to a noticeably lower loss in the second reconstruction. Given the pronounced disparity, if we compute the ratio of the first to the second reconstruction loss, the value tends to be close to 1 for belonging images, but significantly greater than 1 for non-belonging images (see Figure 2(c)).
@@
Motivated by this insight, we propose an image attribution method based on AutoEncoder Double-Reconstruction, namely AEDR. Unlike existing gradient-guided approaches that utilize single reconstruction loss, AEDR employs the loss ratio from double reconstructions, calibrated by an image homogeneity metric. This approach mitigates discrepancies arising from image texture complexity. Additionally, we implement a Kernel Density Estimation method (Kim and Scott 2012) for adaptive threshold selection, which assumes no prior knowledge of underlying data distributions, thereby enhancing the adaptability across diverse generative models.Our contributions are summarized as follows:
@@
  • We reveal that existing reconstruction-based image attribution methods, which rely on the single reconstruction loss, face significant challenges when applied to SOTA generative models. We identify a key discrepancy: under autoencoder-based double-reconstruction, belonging images show significantly higher latent feature consistency across reconstructions than non-belonging ones.
@@
  • We propose AEDR, an training‑free passive attribution method that leverages the ratio of losses from two successive reconstructions and calibrates it based on image homogeneity. AEDR performs direct reconstruction using the model’s autoencoder, eliminating the need for gradient-based optimization or additional training.
@@
  • Extensive experiments on eight SOTA or widely used generative models demonstrate that AEDR improves the attribution accuracy by an average of 25.5% and reduces the attribution time to just 1% of the best baseline.
@@ 2 Related Work
Early frameworks like Generative Adversarial Networks (GANs) (Goodfellow et al. 2014), Variational Autoencoders (VAEs) (Kingma and Welling 2014), and autoregressive methods such as PixelCNN (Van Den Oord et al. 2016) laid the groundwork for realistic image synthesis. However, diffusion models have recently become the leading approach. Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al. 2020) achieve remarkable image fidelity by iteratively adding and reversing Gaussian noise, using simplified objectives and U-Net architectures. Building on this, Latent Diffusion Models (LDMs) improved efficiency by processing diffusion within a compact latent space, enabling rapid, high-resolution image generation with significantly lower computational requirements.
@@
As generative images grow increasingly photorealistic, the need for reliable detection is increasing (Chen et al. 2023). Early work by Marra combined CycleGAN with steganalysis to identify GAN-generated images, laying a foundation for this field. Most existing methods frame the task as a binary classification problem, distinguishing between real and generated images. These approaches often rely on discriminative texture features (Chen and Yashtini 2024; Konstantinidou et al. 2025; Li et al. 2025) or frequency domain signals (Frank et al. 2020; Dong et al. 2022; Corvi et al. 2023) as core detection cues. Recent approaches such as AEROBLADE (Ricker et al. 2024) and HFI (Choi et al. 2024) adopt reconstruction losses via autoencoders and use LPIPS as a detection metric. Although these methods demonstrate promising performance in generative image detection, they do not address the more challenging task of origin attribution.
@@
Current methods for attributing the origin of generated images can be broadly classified into three types: watermark-based methods (Swanson et al. 1996; Wen et al. 2023; Yang et al. 2024, 2025), fingerprint-based methods (Ding et al. 2021; Yu et al. 2019, 2020, 2021), and reconstruction-based methods (Wang et al. 2023, 2024). Watermark-based methods embed model-specific information into the generated image and retrieve this information during detection to identify the origin. Fingerprint-based methods inject unique patterns into the model during training and rely on dedicated classifiers for attribution. However, both active attribution approaches require additional intervention during image generation or model training and may compromise the quality of the generated images. In contrast, reconstruction-based passive methods enable origin attribution without the need for any extra operations. For example, both RONAN (Wang et al. 2023) and LatentTracer (Wang et al. 2024) adopt gradient-based reconstruction approaches, using the loss between the reconstructed and original images as the attribution signal.
Figure 3: The framework of AEDR. Our method consists of three key modules: double-reconstruction based on autoencoder (Section 3.2), calibration mechanism (Section 3.3), and threshold determination via kernel density estimation (Section 3.4). AEDR employs the ratio of double-reconstruction losses, calibrated by an image homogeneity metric, as the attribution signal. The decision threshold is determined via kernel density estimation and applied for origin attribution.
@@ 3 Method
This section provides a detailed description of AEDR, the passive attribution method proposed in this paper. The process begins with two independent reconstructions of the input image using an autoencoder. The ratio of their reconstruction losses is then computed. To improve robustness, this ratio is further calibrated using an image homogeneity metric, yielding the final attribution signal. The attribution outcome is determined by comparing this signal against a predefined threshold. The overall structure of the AEDR framework is depicted in Figure 3.
3.1 Problem Statement and Threat Models
We begin by formalizing the key concepts related to the attribution task, followed by the definition of the attribution goal and the specification of the threat model.
Belonging vs. Non-Belonging Image. Given an image generative model $\mathcal{M}:\mathcal{I}\rightarrow\mathcal{X}_{\mathcal{M}}$ , where $\mathcal{I}$ is the input space and $\mathcal{X}_{\mathcal{M}}\subset\mathcal{X}$ denotes the subset of images produced by $\mathcal{M}$, with $\mathcal{X}$ representing the full image space. A test image $x$ is classified as a belonging image if and only if $x\in\mathcal{X}_{\mathcal{M}}$; if $x\notin\mathcal{X}_{\mathcal{M}}$, it is classified as a non-belonging image.
Attribution Goal. The goal of the auditor is to determine whether a given image $x$ was generated by the target model $\mathcal{M}$, without modifying the training or inference process of $\mathcal{M}$, and without applying any post-processing to its outputs. This is formulated as a binary function $F:\{\mathcal{M},x\}\rightarrow\{0,1\}$, where $F(\mathcal{M},x)=0$ indicates that $x$ is attributed to $\mathcal{M}$, and $F(\mathcal{M},x)=1$ indicates a non-belonging image.
Threat Model. We assume a white-box setting in which the auditor has access only to the autoencoder $\mathcal{R}$ associated with the model $\mathcal{M}$. Specifically, the auditor can query both the encoder and decoder for image reconstruction, but does not have access to model gradients, training data, or internal parameters of $\mathcal{M}$. This is in contrast to prior work, which require full white-box access, including gradient computation. By relying solely on reconstruction queries, AEDR is more practical for deployment in restricted-access scenarios.
3.2 Double-Reconstruction via Autoencoder
The specific workflow is as follows: Given the autoencoder $\mathcal{R}$ of the target model $\mathcal{M}$ and a test image $x$, we perform two successive reconstructions. The two reconstruction steps are respectively defined in Equations (1) and (2).
@@ 3.2 Double-Reconstruction via Autoencoder
Here, $x^{*}$ and $x^{**}$ denote the reconstructed images obtained from the first and second passes through the autoencoder $\mathcal{R}$. The corresponding reconstruction losses are denoted by $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, both computed using mean squared error (MSE). The advantage of using MSE is demonstrated in Experiment 4.5. AEDR focuses on analyzing the trend of variation in reconstruction losses for the test image $x$ during the double-reconstruction process. Specifically, we define an uncalibrated attribution signal $t$ as the ratio between the first and second reconstruction losses, formally given by: